OpenStudy (anonymous):

Find the derivative of the function. f(t) = t^(3/2)*log(base8)(sqrt[t+2]) f'(t)= ?????

ganeshie8 (ganeshie8):

start by converting base_8 to base_e

OpenStudy (zzr0ck3r):

\.( t^{\frac{3}{2}}*\log_8(\sqrt{t+2})\) without the . gives you this \( t^{\frac{3}{2}}*\log_8(\sqrt{t+2})\)

OpenStudy (whpalmer4):

\[\log_b x = \frac{\ln x}{\ln b}\]

OpenStudy (anonymous):

So what about the derivative?